Consider two numbers. The second number is five less than triple the first number. The sum of the two numbers is twenty-three. Find each number.
Answer by LinnW(1048)

Let f = the first number
and
s = the second number
s = -5 + 3f
f+s = 23
solve for f by adding -s to each side
f = 23 - s
Substitute ( 23 - s ) for f in s = -5 + 3f
s = -5 + 3( 23 - s )
s = -5 + 69 -3s
s = 64 -3s
add 3s to each side
4s = 64
divide each side by 4
s = 16
Since s = -5 + 3f, 16 = -5 + 3f
add 5 to each side
21 = 3f
divide each side by 3
7 = f
The first number is 7 and the second is 16.